The angles of elevation of the top of a tower, as seen from two points A and B situated on the same line at distances p and q respectively, from the foot of the tower, are 30 degrees and 60 degrees respectively. Prove that the height of the tower is root(pq).

Grade:10

1 Answers

EPARI SRAGVEE
10 years ago
by drawing the figure, we can get... let height of the tower be....h tan30= h/p =p/root3=h tan60=h/q=qroot3=h on simplification pq/h=h =pq=h2 =h=rootpq